OpenStudy (anonymous):

find the electric field at a distance z above the midpoint of a straight line of length 2l with a uniform charged density d

OpenStudy (anonymous):

well.. can u start?! i would help..

OpenStudy (anonymous):

ya tell me the answer

OpenStudy (anonymous):

Near it, you can use a cylindrical Gaussian surface and ignore ends.

OpenStudy (anonymous):

you need to solve: \[E=\int_{-L}^Lkz\rho dx/(x^2+z^2)^{3/2}\]

OpenStudy (anonymous):

check eq (41): http://integral-table.com/

OpenStudy (anonymous):

I am dividing the rod in to small charges \[dq=\rho dx\] \[dE=kdq/r^2\] you x components cancel each other, and to get y component you need to multiply by cos or sin depending where you chose angle

OpenStudy (anonymous):

\[r=\sqrt{x^2+z^2}\] \[\cos{\theta}=z/\sqrt{x^2+z^2}\]